OpenStudy (unklerhaukus):

2. Express the following as ‘for all’ assertions (using symbols and words): (a) The equation \(x^3 = 28\) does not have a natural number solution. (b) \(0\) is less than every natural number. (c) The natural number n is a prime.

OpenStudy (unklerhaukus):

my answers so far (a) \[(\forall x\in \mathbb N)(x^3\neq28)\] (b) \[(\forall x\in\mathbb N)(0<x)\] (c) \[(\forall n\in \mathbb N)\left((\neg\exists p,q \in \mathbb N)\wedge (pq= n)\right)\]

ganeshie8 (ganeshie8):

last one looks incomplete, i think its not covering (p,q) = (1,n) case

OpenStudy (unklerhaukus):

(c) \[(\forall n\in \mathbb N)\left[((\exists p,q \in \mathbb N)\wedge (pq= n ))\Rightarrow (p=n\vee q=n)\right]\]

ganeshie8 (ganeshie8):

\((\forall (n,p,q)\in \mathbb N)\left[(pq= n )\Rightarrow (p=n\vee q=n)\right] \)

ganeshie8 (ganeshie8):

does it look correct

OpenStudy (unklerhaukus):

your statement looks stronger than mine

OpenStudy (unklerhaukus):

\[\large(\forall n,p,q\in \mathbb N)\left[pq= n \Rightarrow p=n\vee q=n\right]\]

ganeshie8 (ganeshie8):

lol i have no idea if its correct, i just manipulated ur symbols

ganeshie8 (ganeshie8):

i mean the usage of for all, can we pass multiple arguments... ?

OpenStudy (unklerhaukus):

i am not certain

ganeshie8 (ganeshie8):

\(\large(\forall n,p\in \mathbb N)\left[p | n \Rightarrow p\in \{1,n\}\right] \)

ganeshie8 (ganeshie8):

\(\large(\forall n,p\in \mathbb N)\left[p | n \Rightarrow p =1 \vee p=n\right] \)

OpenStudy (unklerhaukus):

what happened to q /

OpenStudy (experimentx):

for all n belongs to natural number, there exist a natural number less than n and not equal to 1 such that n/p does not belong to the set of natural numbers.

OpenStudy (unklerhaukus):

?

OpenStudy (experimentx):

also try separating the latter part with | or : \[ (\forall x\in \mathbb N)|(x^3\neq28) \\ (\forall x\in \mathbb N):(x^3\neq28)\]

OpenStudy (experimentx):

\[ \exists n \in \mathbb N |(\exists p \in\mathbb N )\wedge ( 1 <p < n )\wedge (n/p \cancel{\in} \mathbb N) \]

OpenStudy (unklerhaukus):

im looking for a ‘for all’ assertion

OpenStudy (experimentx):

also if you are looking for just particular prime ... i think there shouldn't be "there exists" before N.

OpenStudy (experimentx):

*n

OpenStudy (anonymous):

for all in this case would mean \(\forall m\in \mathbb{N}, m\not| n\)

OpenStudy (anonymous):

i guess you would have to say also \(m\neq 1\)